What is the oxidation number and the correct formula and working outs for these;

1. Carbon in Sodium Carbonate

2. Sulphur in Hydrogen Sulphide

3. Nitrogen in Potassium Nitrate


1
Expert's answer
2021-05-12T06:25:28-0400

1)Na2CO3=2+x-6=0

X=4

2)H2S=2+X=0

X=-2

3)KNO3=1+X-6=0

X=5